What nominal interest rate compounded semi-annually is equivalent to j4 = 11%?
Answer #1

Let the nominal annual interest rate be x

(1+x/2)(1+x/2) - 1 = 0.11

(1+x/2)^2 = 1.11

X/2 = 5.356%

X = 10.713%

Hence, interest rate of 10.713% compounded semi annually will give 11% effective interest rate per annum.